# Edge-guided Low-light Image Enhancement with Inertial Bregman Alternating Linearized Minimization

### v1

@@ III-C Inertial Bregman Alternating Minimization Algorithm
$\mathcal{P}_{g,\mu}^{\psi}(u;v):=\arg\min_{\hskip-11.38092ptx\in\mathbb{R}^{n}}\left\{g(x)+\langle x,v\rangle+\mu D_{\psi}(x,u)\right\}$ (14)
@@
In the following, we propose an inertial Bregman alternating linearized minimization algorithm to solve the general optimization problem that minimizes the function given in (13). The algorithm framework is summarized in Algorithm 1. In Algorithm 1, the kernel functions $\psi_{k}$ and $\varphi_{k}$, for all $k$, are selected to be strongly convex with the smallest moduli $\rho_{1}$ and $\rho_{2}$, respectively. The most related works are [44, 45, 46, 41, 47].
@@ 1. Compute $y_{1}^{k},z_{1}^{k}$ and $x_{1}^{k+1}$ by
Next, we apply the proposed inertial Bregman alternating minimization algorithm, i.e., Algorithm 1, to solve (12). Note that the model (12) falls into the form of (13) with $(x_{1},x_{2}):=(l,r)$, $\theta_{1}(x_{1}):=\|\nabla l\|_{1}$, $\theta_{2}(x_{2}):=\frac{\alpha}{2}\|\nabla r-\nabla g\|^{2}$, $h^{+}(x_{1},x_{2}):=0$, and $h^{-}(x_{1},x_{2}):=-\frac{\beta}{2}\|l-s-r\|^{2}$. By choosing $\psi_{k}\equiv\psi:=\frac{1}{2}\|\cdot\|^{2}$ and $\varphi_{k}:=\frac{1}{2}\|\cdot\|_{M_{k}}^{2}$ with $M_{k}:=\gamma_{k}I-\frac{\alpha}{\tau_{2}^{k}}\nabla^{\top}\nabla$, where $\gamma_{k}>\frac{\alpha}{\tau_{2}^{k}}\lambda_{\max}(\nabla^{\top}\nabla)$ and applying Algorithm 1 to solve (12), we obtain the following $l$- and $r$-subproblems:
[…]
and
[…]
$r^{k+1}=\frac{\alpha\nabla^{T}\nabla g+\beta(z_{2}^{k}-l^{k+1}-s)+\tau_{2}^{k}(y_{2}^{k}+M_{k}r^{k}-r^{k})}{\gamma^{k}}.$ (19)
[…]
In the implementation, the parameter $\gamma_{k}$ is not difficult to choose since the maximum eigenvalue $\lambda_{\max}(\nabla^{\top}\nabla)\leq 8$ [48].
@@ IV Convergence results
Now we are prepared to establish the convergence of the proposed Algorithm 1. We start by giving some standard assumptions for (13).
@@
    there exist $0<\bar{\alpha}_{i}<\frac{(1-\varepsilon)\rho_{i}}{2}$ such that $\alpha_{i}^{k}\in[0,\bar{\alpha}_{i}]$, $i=1,2$. In addition, $\beta_{i}^{k}\in[0,\bar{\beta}_{i}]$ for some $\bar{\beta}_{i}>0$, $i=1,2$.
@@
    $\delta_{1}=\frac{\bar{\alpha}_{1}+\bar{\beta}_{1}\rho_{1}}{(1-\varepsilon)\rho_{1}-2\bar{\alpha}_{1}}\lambda_{1}^{+},~{}~{}\delta_{2}=\frac{\bar{\alpha}_{2}+\bar{\beta}_{2}\rho_{2}}{(1-\varepsilon)\rho_{2}-2\bar{\alpha}_{2}}\lambda_{2}^{+}.$ (22)
@@
Note that according to Assumption 3, for all $k\in\mathbb{N}$, $\tau_{i}^{k},i=1,2$ is bounded from above, that is, $\tau_{i}^{k}\leq\tau_{i}^{+},i=1,2,$ with
@@
$\Theta_{\delta_{1},\delta_{2}}(\mathbf{w}):=\Phi\left(w_{1}\right)+\frac{\delta_{1}}{2}\left\|w_{11}-w_{21}\right\|^{2}+\frac{\delta_{2}}{2}\left\|w_{12}-w_{22}\right\|^{2},$ (24)
where $\Phi$ is defined in (13), $\delta_{1},\delta_{2}$ are defined in (22), $w_{1}=\left(w_{11},w_{12}\right)\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}},w_{2}=\left(w_{21},w_{22}\right)\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$, and $\mathbf{w}=\left(w_{1},w_{2}\right)$.
@@
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold. Then, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}_{k\in\mathbb{N}}$ is nonincreasing. In particular, for all $k\in\mathbb{N}$ and $\varepsilon>0$, it holds that
@@
The proof of Lemma 1 can be found in the supplementary material Appendix A.
Lemma 2.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold. Then, for any $k\geq 1$, there exists $\gamma>0$ such that
The proof of Lemma 2 can be found in the supplementary material Appendix B.
Theorem 1.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold, and $\{\mathbf{w}^{k}\}$ is assumed to be bounded. Then,
  • (i)
  • (ii)
  • (iii)
The proof of Theorem 1 can be found in the supplementary material Appendix C.
Theorem 2.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold, and $\{\mathbf{w}^{k}\}$ is assumed to be bounded. If $\Phi$ in (13) is a KL function. Then, the sequence $\left\{{\bf w}^{k}\right\}_{k\geq 1}$ has finite length, that is, $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty.$ Hence, the whole sequence $\{{\bf w}^{k}\}_{k\geq 1}$ is convergent.
The proof of Theorem 2 can be found in the supplementary material Appendix D.
V Numerical experiments
To evaluate the effectiveness of the proposed scheme, we first analyze the proposed edge prior with some ablation study. Secondly, the parameters of the proposed model and algorithm are studied, respectively. Then, the real-world low-light images are tested to illustrated the robustness of our method. Furthermore, some state-of-the-art methods are compared to illustrate the superiority of the proposed approach.
Figure 3: Visual comparison of different edge extraction methods. The first column is low-light images; the second column and the last column are the edges obtained by using Eq. (3) and the learned edge by our strategy, respectively.
V-A Implement
In this work, we apply Eq. (3) on 485 training images of the LOL dataset [16] to obtain their corresponding edges. As illustrated in Fig. 2, we extract the edge information from the high-contrast image. Subsequently, we employ this edge feature as the label and train the edge network using the low-contrast input. The implementation of our edge network is carried out using the PyTorch framework, and we update it using the Adam optimizer. The initial learning rate is set to $10^{-4}$ and it is halved every 200 epochs. Once the edge network is learned, we integrate it into the proposed framework to enhance the performance of the image Retinex model.
Figure 4: Study on the number of multi-scale residual block (MSRB) modules. Considering the cost and the performance, we set the number of modules as 6 in our experiment.
Figure 5: Ablation study on edge prior. The images and the corresponding (color) histograms are presented. The first column is the input image, the second column is the result without edge prior, and the last column is the proposed result. Enhanced image with our edge prior has more clear detail.
V-B Effectiveness of edge prior
We propose an edge extraction network, which is comprised of the convolutional layer, multi-scale residual block (MSRB), upsample module, and bottleneck layer. To illustrate the effectiveness of the proposed edge network, Fig. 3 presents the visual comparison of the learned edges and the edges from Eq. (3). For the training, we study the effects of the number of MSRB modules. Theoretically, as the number of MSRB modules increases, the performance of the edge network gets better. In Fig. 4, the loss and PSNR curves show that the learned edges are closer to the labels with 6 MSRB modules. Considering the cost and the performance, in the experiments, we set the number of MSRB as 6 to capture the edge information directly from the input low-light images.
In Fig. 5, we compare the results of the proposed model with and without the edge prior. From the visual results of enhanced images in the first row, the results with the proposed learned edge prior show more detailed information. The second and the last rows display the histogram curves. For the histogram, the x-axis represents the variation in hue, while the y-axis represents the total number of pixels with that specific hue. In general, the left side of the x-axis corresponds to dark areas and the right side corresponds to light areas. Therefore, a histogram of a low-light image will predominantly exhibit data points on the left and center of the plot. In contrast, a histogram of an enhanced image will mostly show data points toward the right and center of the plot. In the last row, the red, green, and blue curves present the distribution of the red, green, and blue color channels, respectively. The histogram curves illustrate the competitive performance of the proposed scheme and the advantages of the proposed edge prior.
Figure 6: Comparison between the alternating minimization algorithm (AMA) and the proposed scheme.
V-C Algorithm and model analysis
In this subsection, we analyze the efficiency of the proposed scheme. First of all, we compare the proposed algorithm with the alternating minimization algorithm (AMA), which is referred to in literature [10]. The energy curves presented in Fig. 6 illustrate that with the same parameter setting, the proposed algorithm is faster and more accurate.
Secondly, the algorithm parameters $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, and $\tau_{1}$, $\tau_{2}$ are studied. According to the Assumption 3, $\alpha_{i}^{k}\in[0,\bar{\alpha}_{i}]$ with $0<\bar{\alpha}_{i}<\frac{(1-\varepsilon)\rho_{i}}{2}$ and $\beta_{i}^{k}\in[0,\bar{\beta}_{i}]$ for some $\bar{\beta}_{i}\in(0,1)$, $i=1,2$. In experiments, we set strongly convex modulus $\rho_{1}$ and $\rho_{2}$ to be $1$, then we have $\alpha_{i}\in[0,0.5)$ and $\beta_{i}\in[0,1)$.
In the experiments, by setting $\bar{\beta}_{i}=1-\epsilon$, $\bar{\alpha}_{i}=\frac{1}{2}-\epsilon$, $\rho_{i}=1$, and $\lambda_{i}^{+}=\beta$, we get $\tau_{i}=\frac{\frac{1+\epsilon}{\epsilon}(2.5+\beta_{i}-2\epsilon)\beta}{1-\alpha_{i}}$. Since $\alpha_{i}$ and $\beta_{i}$ are extrapolation parameters, we first fix them as 00 to find $\tau_{i}$. As analyzed before, $\tau_{i}$ can be decided by $\alpha_{i}$, $\beta_{i}$, and $\beta$. The remaining algorithm parameters are $\tau_{i}$, $i=1,2$, and decided by model parameter $\beta$. Hence, we first study the model parameters $\alpha$ and $\beta$ with fixed $\alpha_{i}$ and $\beta_{i}$. After the optimized $\alpha$ and $\beta$ are obtained, we will focus on the algorithm parameters $\alpha_{i}$ and $\beta_{i}$. We first study $\alpha$ and $\beta$ in interval $[1,200]$ with stepsize $10$. We find that our model is quite robust with parameter $\beta$ while the curve is turbulent alone $\alpha$. Hence we narrow the interval to $[5,50]$ and $[0.1,10]$ use the linespace to find $20$ numbers, respectively. Fig. 7 plots the PSNR and NIQE results along the model parameters $\alpha\in[5,50]$ and $\beta\in[0.1,10]$. The surface indicates that the proposed model is stable to parameters $\alpha$ and $\beta$ in the given interval. Since curves become quite flat as the parameter $\alpha$ from $3$ to $5$ along the $\alpha$ axis, which implies choosing $\alpha$ in the interval $[10,40]$ ensures the best result. Similarly, the interval of model parameter $\beta$ is chosen at $[0.1,5]$.
With the obtained optimized model parameter $\beta$, the algorithm parameter $\tau_{i}$ then controlled by $\alpha_{i}$ and $\beta_{i}$, $i=1,2$. Hence, only $\alpha_{i}\in[0,0.5)$ and $\beta_{i}\in[0,1)$ are studied in the algorithm parameter. For these four algorithm parameters, we choose to tune one parameter in its interval while others are fixed as 00. More specifically, we set the model parameters $\alpha_{i}\in[0.01:0.1:0.5]$ and $\beta_{i}\in[0.01:0.1:1]$. From Fig. 8 and Fig. 9, we know that extrapolation parameters speed up the whole algorithm with convergence.
Figure 7: PSNR ($\uparrow$) and NIQE ($\downarrow$) results along with model parameter $\alpha$ and $\beta$.
V-D Image enhancement
In this section, $15$ low-light images in the dataset RealBlur [49] with paired bright-light images are tested to demonstrate the validity of the proposed approach. Here we indicate these real low-light images as RealBlur15. We compare our method with some state-of-the-art image enhancement models, including the traditional methods (LR3M [12], QSPE [50], DRE [13], RMR [14]) and the deep learning methods (RetinexNet [16], URetinex [18], EFINet [51], RQLLIE [52]).
We present the PSNR and SSIM results of dataset Realblur15 in Table II. The highest and the second highest PSNR and SSIM values are marked as red and blue, respectively. From numerical results, the proposed method with the learned prior is better than the methods with traditional priors in most images. While comparing to the deep learning methods, some results fail to achieve the best one, most of our results still get the second best index. Our method achieves the best averaged overall performance in terms of PSNR and SSIM among all the methods, which shows the effectiveness of our method on real low-light images. Besides the numerical results, the visual results of Img1 from the dataset Realblur15 are presented in Fig 10. The input image is with non-uniformed light and small spark lights, which is the real scene image and is hard to handle. As shown in Fig. 10 (b), (c), and (e), we can see that the small spark lights affect the image enhancement results and make the lights abnormal. Compared to the paired ground-truth image (j), Fig. 10 (b), (d), (f), and (g) have a color bias due to the non-uniformed lights. Although the proposed approach generates the best results among the presented images, the color of (i) is still not exactly the same as the ground-truth image (j). Therefore, image enhancement for real-world images is still a challenging task.
TABLE II: Quantitative results with PSNR and SSIM values on datasets Realblur15. The BEST result is in red whereas the SECOND BEST one is in blue.
Input RetinexNet [16] URetinex [18] EFINet [51] RQLLIE [52] LR3M [12] QSPE [50] DRE [13] RWR [14] Ours
Img1 8.91/0.2466 16.93/0.7078 19.69/0.8665 16.60/0.8372 16.56/0.7948 15.70/0.7926 12.53/0.6727 16.35/0.7780 19.92/0.8498 20.36/0.8723
Img2 8.47/0.2965 14.41/0.6872 25.16/0.8729 15.36/0.7933 13.12/0.6481 14.33/0.7368 11.52/0.6803 15.35/0.7098 19.13/0.8352 18.61/0.8596
Img3 9.49/0.2397 16.82/0.6897 23.76/0.8883 16.34/0.8038 15.36/0.7621 15.98/0.7934 13.22/0.6777 16.75/0.8222 17.85/0.8016 20.33/0.8702
Img4 13.09/0.2390 15.07/0.4498 13.63/0.5800 21.76/0.8239 13.56/0.5715 19.19/0.7236 15.83/0.5378 13.99/0.5171 15.85/0.5985 19.26/0.6801
Img5 10.16/0.2381 15.25/0.7264 17.13/0.8159 16.14/0.8683 17.16/0.8158 18.80/0.8760 13.29/0.6129 14.83/0.8171 14.14/0.7539 19.01/0.8826
Img6 6.71/0.3467 14.01/0.6451 21.61/0.8886 14.82/0.7478 10.72/0.6395 12.20/0.7066 9.22/0.6239 11.71/0.6663 15.05/0.8354 15.98/0.8307
Img7 10.44/0.2075 15.28/0.5945 15.56/0.7664 19.54/0.8835 25.13/0.8608 14.09/0.7214 22.66/0.7505 24.97/0.6755 13.97/0.6918 21.84/0.8677
Img8 8.56/0.2137 17.30/0.7930 18.74/0.8919 16.78/0.8919 26.03/0.9200 14.78/0.8682 19.41/0.5935 23.02/0.8793 11.32/0.5343 22.75/0.9574
Img9 10.38/0.1908 16.05/0.6807 14.82/0.7570 18.56/0.8778 22.70/0.8832 18.23/0.8832 21.62/0.6247 19.20/0.7080 13.91/0.6787 21.51/0.8801
Img10 10.43/0.2186 15.82/0.6982 14.95/0.8126 20.89/0.9346 18.04/0.8316 15.33/0.8019 20.38/0.5057 17.12/0.8373 13.64/0.5881 22.80/0.9131
Img11 10.58/0.2415 14.66/0.5793 14.89/0.7024 18.06/0.8365 21.91/0.8894 15.97/0.8108 16.69/0.7629 18.70/0.7202 17.75/0.8643 19.86/0.8331
Img12 10.90/0.2256 16.69/0.5848 14.17/0.7069 18.64/0.8678 25.52/0.8980 19.39/0.7947 17.51/0.5001 21.27/0.6521 13.74/0.6264 20.54/0.7987
Img13 11.87/0.1596 15.57/0.6384 14.03/0.7377 20.33/0.8734 16.95/0.7662 21.76/0.8336 17.54/0.4619 23.34/0.8198 14.89/0.5067 22.94/0.8586
Img14 9.93/0.2665 14.44/0.5629 18.84/0.7712 17.28/0.8079 17.10/0.7608 16.68/0.7656 14.87/0.6558 17.49/0.7638 20.39/0.8260 18.85/0.7755
Img15 9.30/0.2724 14.86/0.6373 17.67/0.7925 16.29/0.8310 16.05/0.7805 15.99/0.7836 12.36/0.6564 15.42/0.7802 16.25/0.7767 18.55/0.8429
Ave. 9.95/0.2402 15.54/0.6450 17.64/0.7900 17.83/0.8452 18.39/0.7881 16.56/0.7928 15.91/0.6211 17.97/0.7431 15.85/0.71787 20.21/0.8482
(a) Input (10.16/0.2381)
(b) RetinexNet (15.25/0.7264)
(c) EFINet (16.14/0.8683)
(d) RQLLIE (17.16/0.8158)
(e) LR3M (18.80/0.8760)
(f) QSPE (13.29/0.6129)
(g) DRE (14.83/0.8171)
(h) RWR (14.14/0.7539)
(i) Ours (19.01/0.8826)
(j) GT
[…]
TABLE III: Average image enhancement results on non-reference datasets DICM, LIME, and MEF. We highlighted the BEST result in red and the SECOND BEST one in blue.
Datasets Index EFINet Bread RQLLIE LR3M DRE RWR Ours
[…]
(a) Input
(b) EFINet (2.247)
(c) Bread (2.535)
(d) RQLLIE (3.144)
(e) LR3M (3.403)
(f) DRE (2.144)
(g) RWR (3.365)
(h) Ours (2.139)
[…]
(a) Input
(b) EFINet (4.220)
(c) Bread (5.132)
(d) RQLLIE (4.424)
(e) LR3M (4.801)
(f) DRE (4.586)
(g) RWR (5.244)
(h) Ours (3.519)
[…]
(a) Input
(b) EFINet (3.333)
(c) Bread (3.779)
(d) RQLLIE (4.161)
(e) LR3M (3.935)
(f) DRE (3.597)
(g) RWR (4.054)
(h) Ours (3.224)
[…]
V-E Non-reference quality assessment
In this subsection, three classical real-world non-reference datasets, DICM [54], LIME [55], and MEF [56] are tested to verify the robustness of the proposed scheme. The state-of-the-art methods EFINet [51], Bread [53], RQLLIE [52], LR3M [12], DRE [13], RMR [14] are compared. Due to the non-reference of the images, we conduct the non-reference metric ARISM, NIQE, and FADE as our index. Note that the higher the ARISM and the lower the NIQE and FADE, the better the image quality is. In Table. III, the average numerical results of the three datasets are listed. With the help of the learned prior, the proposed method has a better performance compared to the approaches with the traditional priors. For most circumstances, the proposed methods have a competitive result.
In Fig. 11, we present the image enhancement results of the image in dataset DICM. The input image (a) is quite simple to handle compared to other extreme dark images. However, the zoom part of results (c), (d), and (g) remain dark. Compared to other methods, our approach generates the enhanced image (h) with a normal light. An image from dataset LIME with edge feature is presented in Fig. 12. For better understanding, we zoom the texture part, which is a keyboard. Our result (h) has clear details of the keyboard in the enhanced image. From the zoom part of the input image (a), the keyboard is under the non-uniformed light, hence (b), (c), (e), and (f) are with the color bias. For (d) and (g), though the color remains, the details cannot be well preserved. In Fig. 13, we present the image enhancement results of the image in Dataset MEF. Similar color bias happens in (d), (e), and (g). For better visuality, we zoom in on the mug in the image. Due to the help of the proposed edge network, our result (h) has a better enhancement performance while others still have low contrast. In conclusion, the proposed method achieved competitive results in enhancing real-world low-light images, as evidenced by both numerical and visual evaluations. The design of the learned edge prior makes the proposed method particularly effective in handling texture images.
VI Conclusion
This paper proposed an edge-guided Retinex model with a novel inertial Bregman alternating linear algorithm for low-light image enhancement. With the help of deep learning, we trained the edge network with paired low-light images with multi-scale residual blocks to obtain the edge prior. The learned edge prior allows us to directly capture edge features from the low-light image, which helps the image enhancement. Additionally, we designed an efficient inertial Bregman alternating linearized minimization algorithm to tackle the proposed Retinex model. The effectiveness of the proposed algorithm was demonstrated through both theoretical and experimental analysis. To better illustrate the superiority of the proposed scheme, some state-of-the-art models on real-world datasets were compared. Based on the presented experimental results, the proposed method with learned prior and solved by the proposed algorithm achieves competitive performance in enhancing the low-light images. However, we acknowledge that our results still exhibit a color bias issue when dealing with non-uniform real-world low-light images. To address this, we plan to design an efficient model in the future to improve the performance of non-uniformed low-light image enhancement.
Appendix A Proof of Lemma 1
@@
$\left|f(u)-f(v)-\langle u-v,\nabla f(v)\rangle\right|\leq\frac{L_{f}}{2}\|u-v\|^{2}.$
@@
$\leq\theta_{1}\left(x_{1}^{k}\right)+h^{+}\left(x_{1}^{k},x_{2}^{k}\right),$
@@
$\leq\theta_{2}\left(x_{2}^{k}\right)+h^{+}\left(x_{1}^{k+1},x_{2}^{k}\right).$
@@
where $L_{2}^{-}\left(x_{1}^{k+1}\right)$ is the Lipschitz constant of $\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}\right)$ at $x_{2}^{k}$. Summing the above two inequalities and plugging into (A), we get
@@
$+\left\langle x_{1}^{k+1}-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right.$
$+\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},z_{2}^{k}\right)-\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)\right.$
@@
$\quad+\frac{1}{2L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}\left\|\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right\|$
@@
$\rho_{1}\tau_{1}^{k}-L_{1}^{-}\left(x_{2}^{k}\right)-L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}-\alpha_{1}^{k}\tau_{1}^{k}-\delta_{1}=(1+\varepsilon)\delta_{1}-\delta_{1}=\varepsilon\delta_{1},$
@@
Similarly, for the $\tau_{2}^{k}$ and $\delta_{2}$ defined in (21) and (22), it holds that
$\rho_{2}\tau_{2}^{k}-L_{2}^{-}\left(x_{1}^{k+1}\right)-L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}-\alpha_{2}^{k}\tau_{2}^{k}-\delta_{2}=\varepsilon\delta_{2},$

### v2

@@ III-B The proposed Retinex image enhancement scheme
$\mathcal{P}_{g,\mu}^{\psi}(u;v):=\arg\min_{\thinspace x\in\mathbb{R}^{n}}\left\{g(x)+\langle x,v\rangle+\mu D_{\psi}(x,u)\right\}$ (12)
@@
We propose an inertial Bregman alternating linearized minimization algorithm to solve the general optimization problem that minimizes the function given in (11). The algorithm framework is summarized in Algorithm 1, where the kernel functions $\psi_{k}$ and $\varphi_{k}$, for all $k$, are selected to be strongly convex with the smallest moduli $\rho_{1}$ and $\rho_{2}$, respectively.
Remark 1.
Compared with existing methods, Algorithm 1 is novel for solving problem (11), which integrates the general inertial acceleration technique with Bregman regularization. The most closely related works include [45, 46, 47, 42, 48]. However, the inertial technique is not incorporated in [42], and the methods proposed in [45, 46, 47, 42, 48] are tailored to tackle particular instances of problem (11).
Next, we apply the proposed inertial Bregman alternating minimization algorithm, i.e., Algorithm 1, to solve (10). Note that the model (10) falls into the form of (11) with $(x_{1},x_{2}):=(l,r)$, $\theta_{1}(x_{1}):=\Psi(l)$, $\theta_{2}(x_{2}):=\frac{\alpha}{2}\Phi(r,g)$, $h^{+}(x_{1},x_{2}):=0$, and $h^{-}(x_{1},x_{2}):=-\frac{\beta}{2}\|l-s-r\|^{2}$. By choosing $\psi_{k}\equiv\psi:=\frac{1}{2}\|\cdot\|^{2}$ and $\varphi_{k}:=\frac{1}{2}\|\cdot\|_{M_{k}}^{2}$ with $M_{k}:=\gamma_{k}I-\frac{\alpha}{\tau_{2}^{k}}\nabla^{\top}\nabla$, where $\gamma_{k}>\frac{\alpha}{\tau_{2}^{k}}\lambda_{\max}(\nabla^{\top}\nabla)$, we obtain the following $l$- and $r$-subproblems:
[…]
$\hskip-14.45377ptr^{k+1}=\frac{\alpha\nabla^{T}\nabla g+\beta(z_{2}^{k}-l^{k+1}-s)+\tau_{2}^{k}(y_{2}^{k}+M_{k}r^{k}-r^{k})}{\gamma^{k}}.$ (15)
[…]
In the implementation, the parameter $\gamma_{k}$ is not difficult to choose since the maximum eigenvalue $\lambda_{\max}(\nabla^{\top}\nabla)\leq 8$ [49]. Drawing from [16], the image produced by the Retinex model usually tends to be over-enhanced. Consequently, a Gamma correction is incorporated into $\hat{L}=\exp(l^{k+1})$. Thus, the final result is given by $\hat{S}=\hat{L}^{\frac{1}{2.2}}\circ\hat{R}$, where $\hat{R}=\exp(-r^{k+1})$.
@@ IV Convergence results
Now we establish the convergence of Algorithm 1. We start by giving some standard assumptions for (11).
@@
    $\exists\ 0<\bar{\alpha}_{i}<\frac{(1-\varepsilon)\rho_{i}}{2}$ such that $\alpha_{i}^{k}\in[0,\bar{\alpha}_{i}]$, $i=1,2$. In addition, $\beta_{i}^{k}\in[0,\bar{\beta}_{i}]$ for some $\bar{\beta}_{i}>0$, $i=1,2$.
@@
    $\hskip-18.06749pt\delta_{1}=\frac{\bar{\alpha}_{1}+\bar{\beta}_{1}\rho_{1}}{(1-\varepsilon)\rho_{1}-2\bar{\alpha}_{1}}\lambda_{1}^{+},~~\delta_{2}=\frac{\bar{\alpha}_{2}+\bar{\beta}_{2}\rho_{2}}{(1-\varepsilon)\rho_{2}-2\bar{\alpha}_{2}}\lambda_{2}^{+}.$ (20)
@@
According to Assumption 3, for all $k\in\mathbb{N}$, $\tau_{i}^{k},i=1,2$ is bounded from above, that is, $\tau_{i}^{k}\leq\tau_{i}^{+},i=1,2,$ with
@@
$\hskip-7.22743pt\Theta_{\delta_{1},\delta_{2}}(\mathbf{w}):=\Phi\left(w_{1}\right)+\frac{\delta_{1}}{2}\left\|w_{11}-w_{21}\right\|^{2}+\frac{\delta_{2}}{2}\left\|w_{12}-w_{22}\right\|^{2},$ (22)
where $\Phi$ is defined in (11), $w_{1}=\left(w_{11},w_{12}\right)\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}},w_{2}=\left(w_{21},w_{22}\right)\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$, and $\mathbf{w}=\left(w_{1},w_{2}\right)$.
@@
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by Algorithm 1, for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold. Then, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}_{k\in\mathbb{N}}$ is nonincreasing. In particular, for all $k\in\mathbb{N}$ and $\varepsilon>0$, it holds that
@@
$\left|f(u)-f(v)-\langle u-v,\nabla f(v)\rangle\right|\leq\frac{L_{f}}{2}\|u-v\|^{2}.$ (24)
@@
$\hskip-14.45377pt\leq\theta_{1}\left(x_{1}^{k}\right)+h^{+}\left(x_{1}^{k},x_{2}^{k}\right),$
@@
$\hskip-14.45377pt\leq\theta_{2}\left(x_{2}^{k}\right)+h^{+}\left(x_{1}^{k+1},x_{2}^{k}\right).$
@@
where $L_{2}^{-}\left(x_{1}^{k+1}\right)$ is the Lipschitz constant of $\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}\right)$ at $x_{2}^{k}$. Summing the above two inequalities and plugging into (25), we get
@@
$+\left\langle x_{1}^{k+1}-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right\rangle$
$+\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},z_{2}^{k}\right)-\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)\right\rangle$
@@
$\quad+\frac{1}{2L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}\left\|\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right\|^{2}$
@@
$\rho_{1}\tau_{1}^{k}-L_{1}^{-}\left(x_{2}^{k}\right)-L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}-\alpha_{1}^{k}\tau_{1}^{k}-\delta_{1}=(1+\varepsilon)\delta_{1}-\delta_{1}=\varepsilon\delta_{1}$
@@
Similarly, for the $\tau_{2}^{k}$ and $\delta_{2}$ defined in (19) and (20), we have
$\rho_{2}\tau_{2}^{k}-L_{2}^{-}\left(x_{1}^{k+1}\right)-L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}-\alpha_{2}^{k}\tau_{2}^{k}-\delta_{2}=\varepsilon\delta_{2}$